CHANGE THE ORDER OF INTEGRA NON TO EVALUATE sin (x) dx dy
Answer #1

Solution

\mathrm{\int _0^9\:\int _{\sqrt{y}}^3\:sin\left(\pi \:x^3\right)dx\:dy}

\mathrm{\mathrm{Change\:order\:of\:integration}}

\mathrm{\int _0^9\int _{\sqrt{y}}^3\sin \left(\pi x^3\right)dxdy}

\mathrm{\mathrm{Given:}\:0\le \:y\le \:9,\:\sqrt{y}\le \:x\le \:3}

\mathrm{\mathrm{For\:}\sqrt{y}\le \:x\le \:3\mathrm{\:plug\:in\:}y=0\quad \Rightarrow \:0\le \:x\le \:3}

\mathrm{\mathrm{Isolate}\:y\:\mathrm{for}\:\sqrt{y}\le \:x\le \:3\quad \Rightarrow \:0\le \:y\le \:x^2}

\mathrm{=\int _0^3\int _0^{x^2}\sin \left(\pi x^3\right)dydx}

\mathrm{\Rightarrow \int _0^{x^2}\sin \left(\pi x^3\right)dy=\left[\sin \left(x^3\pi \right)y\right]_0^{x^2}=\left[\sin \left(x^3\pi \right)y\right]^{x^2}_0=x^2\sin \left(\pi x^3\right)}

\Rightarrow =\int _0^3x^2\sin \left(\pi x^3\right)dx

\mathrm{\mathrm{Apply\:u-substitution:}\:u=\pi x^3}

\mathrm{\frac{du}{dx}=3\pi x^2}

\mathrm{\Rightarrow \:du=3\pi x^2dx}

\mathrm{\Rightarrow \:dx=\frac{1}{3\pi x^2}du}

\mathrm{=\int \:x^2\sin \left(u\right)\frac{1}{3\pi x^2}du}

\mathrm{=\int \frac{\sin \left(u\right)}{3\pi }du}

Adjust integral boundaries :

\mathrm{x=0\quad \Rightarrow \:u=0}

\mathrm{\:x=3\quad \Rightarrow \:u=27\pi }

\mathrm{=\int _0^{27\pi }\frac{\sin \left(u\right)}{3\pi }du=\frac{1}{3\pi }\cdot \int _0^{27\pi }\sin \left(u\right)du}

\mathrm{\mathrm{Use\:the\:common\:integral}:\quad \int \sin \left(u\right)du=-\cos \left(u\right)}

\mathrm{\mathrm{=\frac{1}{3\pi }\left[-\cos \left(u\right)\right]^{27\pi }_0}=\left[-\cos \left(u\right)\right]^{27\pi }_0=2}

\mathrm{\mathrm{=\frac{1}{3\pi }\cdot \:2}=\frac{2}{3\pi }}

\mathrm{=\frac{2}{3\pi }}